[Nps] [New Logentry] RC Update

jroche at jlab.org jroche at jlab.org
Mon Sep 4 12:20:02 EDT 2023


Logentry Text:
--
The PD estimated this morning that beam will not be available to us until Friday swing shift. Shifts have been cancelled until Friday swing shift. Friday day shifts are on standby. The hall will remain in controlled access until  Friday noon.<p>

Some more details:<p>

Since we last met Friday
<ul>
<li>MCC Hot check out of the Hall C beamline : success
<li>The cryo target was cooled down  to Helium temp.  A leak developed on Saturday so the target had to be warmed up Saturday evening (<a href="https://logbooks.jlab.org/entry/4174124">this hclog</a>). The fix will take a few days. Dave M. estimates :"If all goes well we will be able to condense the target Friday morning for operations starting about 1300 or so Friday."
<li>Simona and Ben R. worked on removing capacitor of the blocks that cannot sustain HV on Saturday. It was a  SUCCESS for the first half they worked on (<a href="https://logbooks.jlab.org/entry/4174111">this hclog</a>). On Monday, they are working on the second half.
<li>30 LED run taken (one per column): working on speeding up the analysis (Casey Morean, Ole, Mark J.). Mitch is going to start the replay for all runs, make a list of the runs and let us know when the replay is completed
<li>Cosmics data with updated HV to balance cosmics amplitude (run 894: original HV but after Simona and Ben’s fix of Saturday, run 897 with updated for cosmics HV)
<li>Punch list of PMTS: not sustaining HV (S and B working at it), not read by DAQ (<a href="https://logbooks.jlab.org/entry/4174128">this hclog</a>), no signal above threshold, PMT reading swapped.  A preliminary list is <a href="https://logbooks.jlab.org/entry/4174228">here</a> . Wassim is working on a more detailed list that he plans to share it at the Analysis meeting today at 4pm.
<li>Anode current and trigger rate display: Bob, Yoapeng and Alex C. are working on these. (<a href="https://logbooks.jlab.org/entry/4174213">this entry</a>)
</ul>

Until we get beam (Friday swing)
<ul>
<li>Finish installing NPS correctors (Dave G., Joe B., Ellen) and measure the field (Bogdan) 
<li>Move NPS to nominal positions (NPS at 9.5 m and 17 deg, HMS already at  16.44 deg)
<li>Fix HMS rotation issue (Steve L.)
<li>Fix cryo target leak (Dave M.)
 </ul>

To be scheduled:
<ul>
<li>TO practical training by Silviu. We need a cold cry target for this.
</ul>


---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4174470
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/nps/attachments/20230904/91c36a26/attachment.html>


More information about the Nps mailing list